# Bulk plasmons in elemental metals

### v1

@@ IV Conclusions
We have presented a thorough study of the frequency and momentum dependent inverse dielectric function of 25 elemental metals, computed from first principles at the RPA level of the many-body theory. Such an approach reproduces well the experimental measurements. The RPA data are used to construct spectral band structures of plasmonic excitations. Such excitations exhibit a complex landscape in inhomogeneous metals, on the basis of the mixed orbital character of the valence states. The plasmonic dispersions are much flatter than expected from the free-electron gas model with an equivalent electronic density or plasmon energy, particularly when most of the spectral weight is distributed among several poles. The energy position of the main MPA pole is used to compute an effective number of electrons and compare with free-electron gas model. The spectral weight of such pole is used as a quantitative measure of the deviation from a single pole PPA dielectric function and provides a general picture of its variability across elemental metals.
We have also introduced MPA($\mathbf{q}$}), a generalized analytical representation of the dynamical part of the inverse dielectric function, as an effective multipole-Padé model depending on both momentum and frequency. With our generalized approach, we have accurately fitted the numerical data, with a limited number of poles ranging from 2 to 15 for all the systems. The obtained MPA($\mathbf{q}$}) representations describe complex features present in the spectral band structures, such as non-parabolic energy and intensity dispersions and constructive and destructive quasiparticle overlapping, leading to band crossings and anti-crossings. We report the specific MPA($\mathbf{q}$}) model for each elemental metal, which can be used as a building block for other works, as a first-principle alternative to over simplified analytical models such as the Drude/Lindhard dielectric function. Moreover, the MPA($\mathbf{q}$}) representation in both momentum and frequency can be fundamental to reduce the computational cost of GWGW/BSE calculations, as a starting point towards GWGW/BSE performed on a $\mathbf{k}$$\mathbf{q}$}-path, which would considerably reduce the computational cost of such calculations.
@@ Acknowledgments
The computations of this work were carried out on UNINETT Sigma2 high performance computing resources (grant NN9650K), and through the project EHPC-EXT-2022E01-022 Joint Undertaking, that granted access to Leonardo-Booster@Cineca, Italy, provided by EuroHPC. Work in Norway is supported by the Research Council of Norway in the MORTY project (664350).
@@
The data generated in this article are openly available.

### v2

@@ IV Conclusions
A comprehensive study of the frequency- and momentum-dependent inverse dielectric function of 26 elemental metals, computed from first principles at the RPA level of theory have been presented. The RPA results are found to accurately reproduce available experimental measurements in the optical limit and were used to evaluate plasma, intra-, and inter-band frequencies, construct spectral band structures of the inverse dielectric function, and develop effective analytical representations of plasmonic excitations, using frequency- and momentum- dependent multipole-Padé approximants, MPA($\mathbf{q}$}).
Plasmonic excitations in inhomogeneous metals were found to exhibit multipole structures, in contrast to the simple free-electron gas assumptions, which arise due to the mixed orbital character of the valence states. The energy position of the main MPA pole in the optical limit, ${\mathbf{q}}\to 0$, is used to compute an effective number of electrons that can be used within the free-electron gas model. In turn, the spectral weight of such pole gives a measure of the unsuitability of the commonly used (single-pole) PPA description. The analysis across all the 26 elemental metals studied, provides a general picture of the limited applicability of PPA.
Moreover, the plasmonic dispersions are much flatter than what would be expected for the corresponding free-electron gas model with an equivalent electronic density or plasma energy. This is particularly the case when the spectral weight is distributed among several poles. Using the MPA($\mathbf{q}$}) scheme introduced here, which is not restricted to free-electron or single-pole descriptions, we have accurately fitted effective MPA($\mathbf{q}$}) models to the RPA numerical data with a relatively small number of poles, ranging from 2 to 16 for all the studied metals. The obtained MPA($\mathbf{q}$}) representations are able to describe the complex features present in the spectral band structures, including non-parabolic energy and intensity dispersions and constructive and destructive quasiparticle overlapping, leading to band crossings and anti-crossings. The reported MPA($\mathbf{q}$}) models can be used as a first-principle alternative to over-simplified analytical models such as the Drude/Lindhard dielectric function, or even as fitting schemes to experimental data. Although the present work focuses on elemental metals at the RPA level, the MPA($\mathbf{q}$}) framework is general in the sense that it is constructed directly from the frequency- and momentum-dependent dielectric response. As such, it can in principle be extended to other classes of materials and levels of theory. The MPA($\mathbf{q}$}) models constructed from RPA data, for example, could be used as a key building block to reduce the computational cost of GWGW/BSE calculations, and as a starting point towards GWGW/BSE performed on a $\mathbf{k}$$\mathbf{q}$}-path, thereby significantly reducing the overall computational cost of such calculations.
@@ Acknowledgments
Work in Norway is supported by the Research Council of Norway in the MORTY project (664350). The computations of this work were carried out on UNINETT Sigma2 high performance computing resources (Grant No. NN9650K), and through Project No. EHPC-EXT-2022E01-022 Joint Undertaking, which granted access to Leonardo-Booster@Cineca, Bologna, Italy, provided by EuroHPC.
@@
The data generated in this article are openly available on Ref. [48].